9. In a loss of weight experiment, what

will be the weight of a cubic metal of

dimensions L = 58 cm on a side and

weight 4450 N in a air when submerged

and suspended by a wire in an open

tank of liquid of density 950 kg/m³


1
Expert's answer
2021-02-05T11:47:31-0500
"W_{apparent}=W-F_B,""W_{apparent}=W-\rho gV_{cube},""W_{apparent}=4450\ N-950\ \dfrac{m}{kg^3}\cdot9.8\ \dfrac{m}{s^2}\cdot(0.58\ m)^3=2633.5\ N."